# Gradient Minimization Using Higher Order MethodsYao Ji and Guanghui Lan

### v1

@@ 5.1 Gradient minimization for convex composite problems
3:   Set $(x_{t},L_{t})=\text{AR}(f,x_{t-1},\tfrac{\sigma_{p+1,t-1}}{3(p+1)9^{p}},L_{t-1}).$
4:   If $\|\nabla f(x_{t})\|>\tfrac{\|\nabla f(x_{t-1})\|}{2},$ then $\sigma_{p+1,t}=\tfrac{\sigma_{p+1,t-1}}{4}.$
@@
Suppose the Assumptions in Theorem 4.1 and $f$ satisfies the uniform convexity with parameter $\sigma_{p+1}(f)$ of order $p+1.$ Then Algorithm 7 computes a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ within at most
$\mathcal{O}\left(\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\left\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\right\rceil\right)$
calls to the $p$-th order oracle.
@@
i) By Theorem 4.1 with $\nu=1,$ for each epoch $t,$ we have
@@
after $\mathcal{O}([{\max\{{pL_{p+1}}(f),L_{0},\theta\}}/{\sigma_{p+1,t-1}}]^{\frac{1}{p+1}})$ calls to the $p$-th order oracle.
ii) Algorithm 7 terminates after
$t\geq T=\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\rceil+\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil$
calls to Algorithm 4. This is because if we have $\|\nabla f(x_{t})\|>{\|\nabla f(x_{t-1})\|}/{2},$ then by (51), we conclude the current $\sigma_{p+1,t-1}$ overestimate the true uniform convexity parameter $\sigma_{p+1}(f).$ Thus we set $\sigma_{p+1,t}={{\sigma_{p+1,t-1}}}/{{4}}.$ However the estimate $\sigma_{p+1,t}$ will no longer change after at most $T_{1}=\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\rceil$ epochs.
On the other hand, if we have $\|\nabla f(x_{t})\|\leq{\|\nabla f(x_{t-1})\|}/{2},$ then by (51), we conclude the current $\sigma_{p+1,t-1}$ is smaller enough, and thus do not need to decrease. Furthermore, $\|\nabla f(x_{t})\|\leq\varepsilon$ after at most $T_{2}=\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil$ epochs.
Combining i), ii), the total calls to the $p$-th order oracle reads
$\textstyle\sum_{t=1}^{T}c_{p}\left[\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1,t-1}}\right]^{\frac{1}{p+1}}=c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}\textstyle\sum_{t=1}^{T}\left(\tfrac{1}{\sigma_{p+1,t-1}}\right)^{\frac{1}{p+1}}$
$\overset{\text{(i)}}{\leq}c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}\textstyle\sum_{t=1}^{T_{1}}\left(\tfrac{4^{t-1}}{\sigma_{p+1,0}}\right)^{\frac{1}{p+1}}+\tfrac{c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}T_{2}}{\sigma_{p+1}(f)^{\frac{1}{p+1}}}$
$\overset{\text{(ii)}}{\leq}c_{p}\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\tfrac{4^{\frac{1}{p+1}}}{4^{\frac{1}{p+1}}-1}+c_{p}\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil,$
where $c_{p}=3(p+1)9^{p},$ and in (i), we used the fact $\sigma_{p+1,t}=\sigma_{p+1,t^{\prime}}$ for all $T_{1}\leq t^{\prime}\leq T,$ and $\sigma_{p+1,T_{1}}=\sigma_{p+1,0}/{4^{T_{1}}}\geq\sigma_{p+1}(f);$ and in (ii) we used the definition of $T_{1}$ and $T_{2}.$
Observe the convergence rate of Theorem 5.3 matches the convergence rate when assuming $\sigma_{p+1}(f)$ is known. Similar guess and check procedure for unknown $\sigma_{p+1}(f)$ can easily extended to when $p+1\neq q,$ and thus omitted for simplicity.
@@ 6 Concluding Remarks
In this paper, we establish a unified framework of high-order accumulative regularization for gradient norm minimization over convex and uniformly convex functions. By carefully designing regularization strategies and leveraging high-order information, the proposed algorithms bridge the long-standing gap between function-residual and gradient-based complexity guarantees. Our results demonstrate that the high order framework effectively accelerates the existing slow convergence of the gradient norm by exploiting the subroutine’s fast function-residual decrease. Furthermore, we develop several parameter-free variants that achieve the same convergence rates without requiring prior knowledge of problem parameters. Overall, our framework attains best-known gradient norm rates across a broad range of smoothness and curvature regimes, recovering classical first-order results as special cases while extending to previously unexplored high-order settings.
@@
We aim to maintain the following relations $\mathcal{R}_{k}^{1},\mathcal{R}_{k}^{2}$ for all $k\geq 1.$ We prove by induction. When $k=1,$ by the definition of $f_{1}(x)$ in Algorithm 1, we have $f_{1}^{*}=f(x_{1})+\sqrt{1/{(L+M)}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}.$ By choosing $A_{1}=1,$ we conclude $\mathcal{R}_{1}^{1}$ holds. Furthermore, we have
$f_{1}(x)$ $\,=f(x_{1})+\tfrac{1}{\sqrt{L+M}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}+\tfrac{C\|x-x_{0}\|^{3}}{6}$
$\overset{\text{(a)}}{\leq}\min\limits_{y\in\mathbb{R}^{n}}\left[f(y)+\tfrac{2L\|y-x_{0}\|^{3}}{6}\right]+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}+\tfrac{C\|x-x_{0}\|^{3}}{6}$
$\,\leq f(x)+\tfrac{(2L+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}},$
where in (a), we used [30, Lemma 4], [23, Lemma 6], and in the last step, we choose $y=x.$ Thus, $\mathcal{R}_{1}^{2}$ holds. Assume that $\mathcal{R}_{k}^{1}$ and $\mathcal{R}_{k}^{2}$ hold for for some $k\geq 1.$ For $k+1,$
@@
$\overset{\mathcal{R}_{k}^{2}}{\leq}A_{k}f(x)+\tfrac{(2L+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}$
@@
$\overset{\text{(b)}}{\leq}A_{k+1}f(x)+\tfrac{(2L+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}},$
where in (b), we used the convexity of $f$ and $A_{k+1}=A_{k}+a_{k}.$ Therefore $\mathcal{R}_{k+1}^{2}$ holds. It remains to show $\mathcal{R}_{k+1}^{1}$ holds, which it is a straightforward modification of [23], thus omitted for simplicity. By the definition of $f_{k+1}(x),$ we have
@@
We proceed with providing a lower bound on $f_{k}(x).$ By the definition of $f_{k}(x)$ in Algorithm 1, we have $f_{k}(x)=\ell(x)+\textstyle\sum_{i=1}^{k-1}a_{i}\left(f(x_{i+1})+\langle\nabla f(x_{i+1}),x-x_{i+1}\rangle\right)+{C}\|x-x_{0}\|^{3}/6,$ therefore, $f_{k}(x)$ is an uniform strongly convex function. Thus, we have
(53) $f_{k}(x)\overset{\text{(b)}}{\geq}$ $f_{k}^{*}+\tfrac{C}{12}\|x-\nu_{k}\|^{3}\ \overset{\mathcal{R}_{k}^{1}}{\geq}A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}{A_{j}\tfrac{\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}}+\tfrac{C\|x-\nu_{k}\|^{3}}{12},$
where in (b), we used the uniform strong convexity of $f_{k}$ $\nu_{k}$ is the solution of $f_{k}(x)$ defined in step 2 of Algorithm 1, and $\nabla f_{k}(\nu_{k})=0.$ Substituting the lower bound for $f_{k}(x)$ into (52), we have
@@
$\overset{\text{(a)}}{\geq}A_{k+1}[f(x_{k+1})+\langle\nabla f(x_{k+1}),y_{k}-x_{k+1}\rangle]+\textstyle\sum_{j=1}^{k}{\tfrac{A_{j}\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}}$
@@
where in (a), we used the convexity of $f$ and $A_{k+1}=A_{k}+a_{k}.$ By [30, Lemma 4], [23, Lemma 6], if $M\geq 2L,$ there holds
(55) $\langle\nabla f(x_{k+1}),y_{k}-x_{k+1}\rangle\geq\sqrt{\tfrac{2}{L+M}}\|\nabla f(x_{k+1})\|^{\frac{3}{2}}.$
@@
$A_{k}$ $=A_{k-1}+a_{k}=\tfrac{k(k+1)(k+2)}{6},\quad a_{k}^{-\frac{3}{2}}A_{k+1}\geq\tfrac{2}{3},\quad\tfrac{1}{L+M}\geq\tfrac{4}{C(\sqrt{2}-1)^{2}}.$
@@
$\tfrac{\left(\sqrt{2}-1\right)A_{k+1}\left\|\nabla f(x_{k+1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}{\geq}\tfrac{4a_{k}^{\frac{3}{2}}\left\|\nabla f(x_{k+1})\right\|^{\frac{3}{2}}}{3\sqrt{C}}$
@@
$A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}A_{j}\tfrac{\|\nabla f(x_{j})\|^{\frac{3}{2}}}{\sqrt{L+M}}\overset{\text{$\mathcal{R}_{k}^{1}$}}{\leq}f_{k}^{*}\leq f_{k}(x)\overset{\text{$\mathcal{R}_{k}^{2}$}}{}$
${\leq}A_{k}f(x)+\left[\left(\tfrac{1}{2\sqrt{2}}+\tfrac{1}{6}\right)(L+M)+\tfrac{C}{6}\right]\|x_{0}-x\|^{3}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L+M}}.$
@@
For a general $p$-th order accelerated tensor method, we can utilize [25, Corollary 1] and analyze a modified accelerated tensor method to derive a slow gradient complexity $\mathcal{O}(1/\varepsilon^{p})$, which satisfies 1 and thus can be used as a subroutine in the high-order AR framework. The proof follows the same structure as in [25] by modifying the estimating sequence and is analogous to the accelerated CNM in how the gradient-norm output is obtained; therefore, we omit the details for simplicity.

### v2

@@ 5.2 Parameter-free gradient minimization for uniformly convex functions
[…]
4:   If $\|\nabla f(x_{t})\|>{\|\nabla f(x_{t-1})\|}/{2},$ then $\sigma_{p+1,t}={\sigma_{p+1,t-1}}/{4}.$
@@
Suppose the assumptions in Theorem 4.1 hold and that $f$ is uniformly convex of order $p+1$ with parameter $\sigma_{p+1}(f)$. Then Algorithm 7 computes a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ within at most
$\mathcal{O}\left(\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\left\{\left\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\right\rceil+\left\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\right\rceil\right\}\right)$
calls to the $p$-th-order oracle, where $\theta$ is a user-defined inexactness parameter in 3.
@@
i) By Theorem 4.1 with $\nu=1$ and uniform convexity with $q=p+1,$ for each epoch $t,$ we have
@@
after $\mathcal{O}([{\max\{{pL_{p+1}}(f),L_{0},\theta\}}/{\sigma_{p+1,t-1}}]^{\frac{1}{p+1}})$ calls to the $p$-th-order oracle.
ii) At epoch $t,$ if $\|\nabla f(x_{t})\|>{\|\nabla f(x_{t-1})\|}/{2},$ then by (52), we conclude the current $\sigma_{p+1,t-1}$ overestimates the true uniform convexity parameter $\sigma_{p+1}(f).$ Hence, we set $\sigma_{p+1,t}={{\sigma_{p+1,t-1}}}/{{4}}.$ However, the estimate $\sigma_{p+1,t}$ will no longer change after at most $\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\rceil$ epochs. Denote the set of such epoch indices by $\mathcal{T}_{1}.$ On the other hand, at epoch $t,$ if $\|\nabla f(x_{t})\|\leq{\|\nabla f(x_{t-1})\|}/{2},$ then by (52), we conclude the current $\sigma_{p+1,t-1}$ is small enough, and thus in the next epoch $t,$ we have $\sigma_{p+1,t}=\sigma_{p+1,t-1}.$ Furthermore, $\|\nabla f(x_{t})\|\leq\varepsilon$ after at most $\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil$ epochs. Denote the set of such epoch indices by $\mathcal{T}_{2}.$ Therefore, Algorithm 7 terminates in at most
$\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\rceil+\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\rceil$
epochs of the parameter-free AR method Algorithm 4. Denote $\mathcal{T}\coloneqq\mathcal{T}_{1}\cup\mathcal{T}_{2}.$ Combining i), ii), we conclude the total number of $p$-th-order oracle calls is
$\textstyle\sum_{t\in\mathcal{T}}c_{p}\left[\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1,t-1}}\right]^{\frac{1}{p+1}}=c_{p}\max\{{pL_{p+1}}(f),L_{0},\theta\}^{\frac{1}{p+1}}\textstyle\sum_{t\in\mathcal{T}}\left(\tfrac{1}{\sigma_{p+1,t-1}}\right)^{\frac{1}{p+1}}$
$\overset{\text{(i)}}{\leq}c_{p}\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}|\mathcal{T}_{1}|+c_{p}\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)/4}\right)^{\frac{1}{p+1}}|\mathcal{T}_{2}|,$
where $c_{p}=3(p+1)9^{p},$ and in (i), we used the fact $\sigma_{p+1,t}>\sigma_{p+1}(f),$ for all $t\in\mathcal{T}_{1},$ and $\sigma_{p+1,t-1}=\sigma_{p+1,t}$ for all $t\in\mathcal{T}_{2}.$ Suppose $t_{0}$ is the smallest index such that $\sigma_{p+1,t_{0}}=\sigma_{p+1,t},$ then by (52) we conclude that $\sigma_{p+1}(f)\leq\sigma_{p+1,t_{0}-1},$ and $\sigma_{p+1,t_{0}-1}=4\sigma_{p+1,t_{0}}.$ Therefore, for all $t\in\mathcal{T}_{2},$ we have $\sigma_{p+1,t}\geq\sigma_{p+1}(f)/4.$ Substituting the definitions of $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$ concludes the proof.
Observe that the rate matches the case where $L_{p+1}(f),\sigma_{p+1}(f)$ are known. Similar guess-and-check procedures for unknown $\sigma_{p+1}(f)$ can be extended to the case $p+1\neq q,$ and are thus omitted for brevity.
@@ 6 Concluding Remarks
In this paper, we establish a unified high-order accumulative regularization framework for gradient-norm minimization on convex and uniformly convex objectives. By carefully designing regularization strategies and leveraging high-order information, the proposed algorithms bridge the long-standing gap between function-residual and gradient norm complexity guarantees. Our results demonstrate that the high-order framework effectively accelerates the existing slow convergence of the gradient norm by exploiting the fast function-residual convergences of the subroutine. Furthermore, we develop several parameter-free variants that achieve the same convergence rates without requiring prior knowledge of problem parameters. Overall, our framework attains the best-known gradient norm rates across a broad range of smoothness and curvature regimes, recovering classical first-order results as special cases while extending to previously unexplored high-order settings.
@@
We aim to maintain relations $\mathcal{R}_{k}^{1},\mathcal{R}_{k}^{2}$ for all $k\geq 1.$ We prove by induction. When $k=1,$ by the definition of $f_{1}(x)$ in Algorithm 1, we have $f_{1}^{*}=f(x_{1})+\sqrt{1/{(L_{3}(f)+M)}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}.$ By choosing $A_{1}=1,$ we conclude $\mathcal{R}_{1}^{1}$ holds. Furthermore, we have
$f_{1}(x)$ $\,=f(x_{1})+\tfrac{1}{\sqrt{L_{3}(f)+M}}\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}+\tfrac{C\|x-x_{0}\|^{3}}{6}$
$\overset{\text{(a)}}{\leq}\min\limits_{y\in\mathbb{R}^{n}}\left[f(y)+\tfrac{2L_{3}(f)\|y-x_{0}\|^{3}}{6}\right]+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}+\tfrac{C\|x-x_{0}\|^{3}}{6}$
$\,\leq f(x)+\tfrac{(2L_{3}(f)+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}},$
where in (a), we used [nesterov2006cubic, Lemma 4], [nesterov2008accelerating, Lemma 6]. Thus, $\mathcal{R}_{1}^{2}$ holds. Assume that $\mathcal{R}_{k}^{1}$ and $\mathcal{R}_{k}^{2}$ hold for some $k\geq 1.$ For $k+1,$ we have
@@
$\overset{\mathcal{R}_{k}^{2}}{\leq}A_{k}f(x)+\tfrac{(2L_{3}(f)+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}$
@@
$\overset{\text{(b)}}{\leq}A_{k+1}f(x)+\tfrac{(2L_{3}(f)+C)\|x-x_{0}\|^{3}}{6}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}},$
where in (b), we used the convexity of $f$ and $A_{k+1}=A_{k}+a_{k}.$ Therefore $\mathcal{R}_{k+1}^{2}$ holds. It remains to show $\mathcal{R}_{k+1}^{1}$ holds, which it is a straightforward modification of [nesterov2008accelerating], we include it here for completeness. By the definition of $f_{k+1}(x),$ we have
@@
We proceed with providing a lower bound on $f_{k}(x).$ By the definition of $f_{k}(x)$ in Algorithm 1, we have $f_{k}(x)=\ell(x)+\textstyle\sum_{i=1}^{k-1}a_{i}\left(f(x_{i+1})+\langle\nabla f(x_{i+1}),x-x_{i+1}\rangle\right)+{C}\|x-x_{0}\|^{3}/6,$ therefore, $f_{k}(x)$ is a uniform convex function of degree 3 (modulus $C/12$). Thus, we have
(54) $f_{k}(x)\overset{\text{(c)}}{\geq}$ $f_{k}^{*}+\tfrac{C}{12}\|x-\nu_{k}\|^{3}\ \overset{\mathcal{R}_{k}^{1}}{\geq}A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}{A_{j}\tfrac{\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}}+\tfrac{C\|x-\nu_{k}\|^{3}}{12},$
where in (c), we used the uniform strong convexity of $f_{k}.$ Note that $\nu_{k}$ minimizes $f_{k}(x)$ defined in Algorithm 1, and $\nabla f_{k}(\nu_{k})=0.$ Substituting the lower bound for $f_{k}(x)$ into (53), we have
@@
$\overset{\text{(d)}}{\geq}A_{k+1}[f(x_{k+1})+\langle\nabla f(x_{k+1}),y_{k}-x_{k+1}\rangle]+\textstyle\sum_{j=1}^{k}{\tfrac{A_{j}\left\|\nabla f(x_{j})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}}$
@@
where in (d), we used the convexity of $f$ and $A_{k+1}=A_{k}+a_{k}.$ By [nesterov2006cubic, Lemma 4], [nesterov2008accelerating, Lemma 6], if $M\geq 2L_{3}(f),$ we have
(56) $\langle\nabla f(x_{k+1}),y_{k}-x_{k+1}\rangle\geq\sqrt{\tfrac{2}{L_{3}(f)+M}}\|\nabla f(x_{k+1})\|^{\frac{3}{2}}.$
@@
$A_{k}$ $=A_{k-1}+a_{k}=\tfrac{k(k+1)(k+2)}{6},\quad a_{k}^{-\frac{3}{2}}A_{k+1}\geq\tfrac{2}{3},\quad\tfrac{1}{L_{3}(f)+M}\geq\tfrac{4}{C(\sqrt{2}-1)^{2}}.$
@@
$\tfrac{\left(\sqrt{2}-1\right)A_{k+1}\left\|\nabla f(x_{k+1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}{\geq}\tfrac{4a_{k}^{\frac{3}{2}}\left\|\nabla f(x_{k+1})\right\|^{\frac{3}{2}}}{3\sqrt{C}}$
@@
$A_{k}f(x_{k})+\textstyle\sum_{j=1}^{k}A_{j}\tfrac{\|\nabla f(x_{j})\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}\overset{\text{$\mathcal{R}_{k}^{1}$}}{\leq}f_{k}^{*}\leq f_{k}(x)$
$\overset{\text{$\mathcal{R}_{k}^{2}$}}{\leq}A_{k}f(x)+\left[\left(\tfrac{1}{2\sqrt{2}}+\tfrac{1}{6}\right)(L_{3}(f)+M)+\tfrac{C}{6}\right]\|x_{0}-x\|^{3}+\tfrac{\left\|\nabla f(x_{1})\right\|^{\frac{3}{2}}}{\sqrt{L_{3}(f)+M}}.$
@@
For a general $p$-th-order accelerated tensor method, we can utilize [nesterov2021implementable, Corollary 1] and analyze a modified accelerated tensor method to derive a slow gradient complexity $\mathcal{O}(1/\varepsilon^{p})$, which satisfies 1 and thus can be used as a subroutine in the high-order AR framework Algorithm 3. The proof follows a similar idea to [nesterov2021implementable], modifying the estimating sequence; it is also analogous to the accelerated CNM in how the gradient norm output is obtained, so we omit the details.